Factory Overhead Rates, Entries, and Account Balance

Eclipse Solar Company operates two factories. The company applies factory overhead to jobs on the basis of machine hours in Factory 1 and on the basis of direct labor hours in Factory 2. Estimated factory overhead costs, direct labor hours, and machine hours are as follows:

Factory 1 Factory 2
Estimated factory overhead cost for fiscal
year beginning August 1 $18,500,000 $44,000,000
Estimated direct labor hours for year 800,000
Estimated machine hours for year 1,250,000
Actual factory overhead costs for August $1,515,800 $3,606,300
Actual direct labor hours for August 64,500
Actual machine hours for August 105,000

a. Determine the factory overhead rate for Factory 1. Round your answer to two decimal places.
$ per machine hour

b. Determine the factory overhead rate for Factory 2.
$ per direct labor hour

Feedback

c. Journalize the entries to apply factory overhead to production in each factory for August. If an amount box does not require an entry, leave it blank.

Factory 1
Factory 2        

Feedback

d. Determine the balances of the factory overhead accounts for each factory as of August 31, and indicate whether the amounts represent overapplied factory overhead or underapplied factory overhead.

Factory 1 $
Factory 2 $

Feedback

Feedback
Answer #1
a
Factory overhead rate for Factory 1 14.80 per machine hour =18500000/1250000
b
Factory overhead rate for Factory 2 55.00 per direct labor hour =44000000/800000
c
Factory 1
Work in Process 1554000 =105000*14.80
         Factory Overhead 1554000
Factory 2
Work in Process 3547500 =64500*55
         Factory Overhead 3547500
d
Factory 1 38200 Credit Overapplied =1554000-1515800
Factory 2 58800 Debit Underapplied =3606300-3547500